# MedFormer: Hierarchical Medical Vision Transformer with Content-Aware Dual Sparse Selection Attention

### v1

@@ 4.2.1 Settings
The classification performance of MedFormer is evaluated on three medical image datasets. To ensure a fair and comprehensive evaluation, MedFormer is compared with two types of methods: specialized medical image classification methods and widely-used Transformer backbone networks. The AdamW optimizer is used with a weight decay of 0.05 and the initial learning rate is set to $1\times 10^{-6}$. A linear warm-up strategy is employed to gradually increase the learning rate during the first 5 epochs, followed by a cosine annealing schedule with a minimum learning rate of $1\times 10^{-5}$. To further enhance the model’s generalization ability, a series of data augmentation techniques are applied, including mixup, cutmix, and random erasing. Additionally, the input images are resized to $224\times 224$, and the training is conducted for 300 epochs. For MedFormer, key parameters are configured, such as setting the region partition factor $S$ to 7, the dimension of each attention head to 32, and the MLP expansion ratio to 3.
@@ 4.2.2 Results
All classification results are summarized in Table 3. It is observed that MedFormer-B achieves the best performance across all three classification datasets. Notably, MedFormer-T achieves comparable performance to some methods with only 3.2M parameters and 0.58G FLOPs. In contrast, HiFuse-Tiny requires significantly more resources but performs well only on the ColonPath dataset. KAT and MG-Trans are effective in reducing computational requirements, but their use of kernel techniques or graph transformers may overlook the extraction of detailed features, thus limiting the improvement in accuracy. The accuracy of MedFormer surpasses that of the compared backbone networks, demonstrating MedFormer’s superior feature representation ability. Additionally, MedFormer with pre-trained weights achieves an accuracy of 100% and 99.92% on the ColonPath and Brain Tumor datasets, respectively, showcasing its immense potential in medical image classification tasks.
Table 6: Results of different methods for Medical Image Detection on Kvasir-Seg and Brain Tumor Detection dataset.
Methods Kvasir-Seg Brain Tumor Detection
$mAP$ $AP_{50}$ $AP_{75}$ Recall Params (M) FLOPs (G) $mAP$ $AP_{50}$ $AP_{75}$ Recall Params (M) FLOPs (G)
RetinaNet* [46] 51.3 87.6 57.7 92.3 36.33 13.05 35.9 70.1 32.6 82.4 36.37 13.11
Faster R-CNN* [47] 48.9 85.4 57.7 90.8 41.35 26.21 37.9 72.1 35.8 77.7 41.36 26.23
Deformable DETR* [48] 54.4 86.1 62.5 96.2 41.07 15.21 38.5 73.4 37.1 87.0 41.08 15.21
PVT-S*[19] 56.2 84.4 62.0 90.0 32.58 12.51 45.7 76.7 47.9 83.7 32.62 12.56
Swin-T*[15] 55.9 89.1 62.5 95.1 36.82 14.58 41.9 75.3 41.2 86.0 36.86 14.64
DAT-T*[26] 47.3 85.2 49.5 92.3 32.05 12.83 41.5 74.4 41.3 85.9 32.09 12.89
BiFormer-S*[16] 49.1 88.4 48.8 95.2 33.62 13.32 45.5 76.9 47.7 85.6 33.66 13.38
MedFormer-T* 52.7 88.9 58.1 95.4 11.09 8.17 42.7 74.8 43.5 86.3 11.13 8.23
MedFormer-S* 53.8 89.3 60.7 93.1 21.22 10.29 44.0 76.1 45.3 84.9 21.27 10.35
MedFormer-B* 56.4 90.1 62.7 94.6 37.59 13.34 45.9 77.2 48.2 84.3 37.63 13.80
@@ 4.3.1 Settings
Under the widely-used UPerNet segmentation framework [49], the performance of MedFormer and popular backbone networks in segmentation tasks is evaluated. Furthermore, comparisons are conducted between MedFormer and several state-of-the-art medical segmentation methods to emphasize its competitiveness. For the ISIC-2018 Segmentation and CVC-ClinicDB datasets, the input image size is set to $256\times 256$, and the Adam optimizer is used with an initial learning rate of $41e-$ 4 for 200 epochs of training. For the Synapse dataset, the input size is adjusted to $224\times 224$, the AdamW optimizer is utilized, and training runs for 400 epochs using a cosine annealing learning rate schedule with an initial learning rate of $31e-$ 3. To avoid overfitting, data augmentation techniques such as horizontal flipping, vertical flipping, and scaling and rotation are applied. For the key parameters of MedFormer, the dropout rate is set to 0.2, while the region partition factor $S$ is 8 for a resolution of $256\times 256$ and 7 for $224\times 224$.
@@ 4.3.2 Results
Table 4 presents the experimental results on the Synapse dataset. MedFormer-B achieves the best average DSC of 84.07%, indicating the significant advantage of MedFormer in accurately overlapping the predicted results with the Ground Truth. However, MedFormer does not achieve the highest score in the HD metric, showing a deficiency in learning edge information. Furthermore, the segmentation results on the ISIC-2018 Segmentation and CVC-ClinicDB dataset are shown in Table 5. For the ISIC-2018 Segmentation dataset, MedFormer outperforms most comparison methods, with only a marginal performance gap compared to BRAU-Net++. However, this slight advantage of BRAU-Net++ comes at the cost of significantly higher computational complexity. MedFormer achieves competitive results by utilizing a simpler segmentation head. For the CVC-ClinicDB dataset, MedFormer achieves the best performance across all metrics, further validating its effectiveness in complex segmentation tasks. In summary, MedFormer demonstrates robust performance on various segmentation tasks, showcasing its potential as a versatile backbone network.
@@ 4.4.1 Settings
Since Transformer-based medical image detection methods with available code are scarce, we reproducing several milestone object detection approaches to highlight the competitiveness of MedFormer in detection tasks. Furthermore, we also compare MedFormer with several popular backbone networks under the RetinaNet[46] detection framework to verify its effectiveness. The input resolution is set to $256\times 256$, and the AdamW optimizer is employed with a cosine annealing learning rate scheduler and a weight decay of $41e-$ 4. The initial learning rate is set to $41e-$ 4. The parameters for MedFormer are adjusted as follows: the region partition factor $S$ is set to 8, and the drop rate is set to 0.3.
@@ 4.4.2 Results
We use four metrics to evaluate the detection performance of all methods: mean Average Precision ($mAP$), $AP_{50}$, $AP_{75}$, and Recall. The detailed results are presented in Table 6. MedFormer-B achieves the best results on the Kvasir-Seg and Brain Tumor Detection datasets in terms of three metrics: $mAP$, $AP_{50}$ and $AP_{75}$. This proves the strong object detection performance of MedFormer. Furthermore, we observe that the differences in detection results among the comparative methods on the Kvasir-Seg dataset are not significant. However, on the Brain Tumor Detection dataset, dedicated detection methods generally lag behind the backbone networks. This could be because these meticulously designed backbone networks possess stronger feature representation capabilities. From these results, we can conclude that MedFormer achieves a remarkable balance between efficiency and accuracy, highlighting its potential as a versatile backbone for medical image recognition.
@@
Dual Sparse Selection 88.08 79.44
Table 8: Ablation study of the number of $k_{1}$ and $\lambda$ on ISIC-2018 Challenge Task 3 dataset.
$k_{1}$ $\lambda$ FLOPs (G) Acc (%)
1,2,8,49 1/8 2.15 87.05
1,4,16,49 1 2.23 87.44
1/4 2.18 87.56
1/8 2.18 88.08
1/16 2.17 85.84
2,8,32,49 1/8 2.25 87.56
@@ 4.5.1 The effectiveness of DSSA
4.5.2 Choices of $k_{1}$ and $\lambda$
In this work, we propose two key parameters, $k_{1}$ and $\lambda$, which control the sparse selection at the region-level and pixel-level, respectively. We gradually increase $k_{1}$ at each stage to ensure sufficient feature extraction in smaller regions. Table 8 presents the experimental results for the combination of these parameters. Based on the analysis of $k_{1}$ values, we observe that the number of tokens each query can attend to may be limited, introducing too many tokens actually lead to performance degradation. Additionally, we test three different $\lambda$ values (the case of $\lambda=1$ is discussed in detail in the next section) and find that the intermediate $\lambda=1/8$ yields the best performance. We believe this outcome is due to the model requiring an appropriate number of pixel-level tokens to effectively learn features. Retaining too many pixel-level tokens may introduce excessive noise, distracting the model from key features, while retaining too few may result in insufficient information, limiting the generalization ability of the model.
Figure 4: Visualization of attention maps from sample images in the ISIC 2018 Classification dataset (top) and the CVC-ClinicDB segmentation dataset (bottom), illustrating the effects of excluding noise tokens. It can be seen that pixel-level sparse selection enhances focus on lesion areas, which is particularly important for extracting critical features.
Figure 5: Visualization of attention maps on ISIC-2018 Classification dataset. For each input, we select two different query positions to observe the region-level tokens they attend to and the pixel-level tokens within those regions. The heatmap values of the pixel-level tokens vary based on their relevance to the query.
@@ 4.5.3 Effectiveness of Excluding Noise
We design ablation experiments to confirm that pixel-level sparse selection can effectively exclude noise. First, we present qualitative results in Table 8. The accuracy of the model significantly decrease when the selection is not applied ($\lambda=1$). We attribute this drop to the introduction of excessive noise tokens, which prevent the model from extracting relevant features. Second, we visualize the attention map with and without pixel-level sparse selection in Fig. 4. The results clearly show that excluding noise tokens allows the model to focus more on the lesion area.
@@ 4.6.1 Visualization of Attention Map
To intuitively demonstrate the superiority of the proposed DSSA, we visualize the attention maps of DSSA at both the region and pixel levels, as shown in Fig. 5. We observe that DSSA first selects the most relevant regions and then focuses on the tokens most related to the query within those regions. This aligns well with our design philosophy and validates the rationality and effectiveness of our approach. In addition, we visualize the regional attention changes across each layer in MedFormer, as shown in Fig. 6. It is clear that the attention regions of the same query vary significantly across different layers. Therefore, applying our proposed dynamic DSSA mechanism at each layer is essential for capturing the most relevant features to the query and for improving the model’s performance.
@@ 4.6.2 Visualization of Semantic Segmentation
We visualize the segmentation results of MedFormer, state-of-the-art medical image segmentation methods and popular backbone networks on three datasets. Fig. 7 shows the results on the Synapse dataset, where MedFormer’s segmentation results closely matches the ground truth. In contrast, other methods exhibit issues such as class misidentification and loss of internal details. Fig. 8 presents the results on the ISIC-2018 Segmentation and CVC-ClinicDB datasets. For ISIC-2018 Segmentation, all methods produce results that are close to the ground truth. However, on the CVC-ClinicDB dataset, some methods fail to accurately segment the lesion boundaries. Overall, MedFormer demonstrates superior performance in segmentation tasks due to its unique dual sparse selection attention mechanism, which enables precise extraction of detailed features.
Figure 8: Visualizations of Segmentation using various methods on the CVC_ClinicDB and ISIC-2018 Segmentation datasets. The lesion regions of GT are shown in red, while the segmentation results of methods are shown in pink.

### v2

@@ 4.2.1 Settings
The classification performance of MedFormer is evaluated on three medical image datasets. To ensure a fair and comprehensive evaluation, MedFormer is compared with two types of methods: specialized medical image classification methods and widely-used Transformer backbone networks. The AdamW optimizer is used with a weight decay of 0.05 and the initial learning rate is set to 1e61e-6. A linear warm-up strategy is employed to gradually increase the learning rate during the first 5 epochs, followed by a cosine annealing schedule with a minimum learning rate of 1e51e-5. To further enhance the model’s generalization ability, a series of data augmentation techniques are applied, including mixup, cutmix, and random erasing. Additionally, the input images are resized to $\times224$, and the training is conducted for 300 epochs. For MedFormer, key parameters are configured, such as setting the region partition factor $S$ to 7, the hyperparameter $\lambda$ to 1/8, the dimension of each attention head to 32, and the MLP expansion ratio to 3.
Table 6: Results of different methods for Medical Image Detection on Kvasir-Seg and Brain Tumor Detection dataset. * indicates the model is initialized with pre-trained weights.
Methods Kvasir-Seg Brain Tumor Detection
mAPmAP $AP_{50}$ $AP_{75}$ Recall Params (M) FLOPs (G) mAPmAP $AP_{50}$ $AP_{75}$ Recall Params (M) FLOPs (G)
RetinaNet* [48] 51.2 87.5 57.4 91.9 36.33 13.05 35.8 70.1 32.5 82.3 36.37 13.11
Faster R-CNN* [49] 48.7 85.3 57.5 90.7 41.35 26.21 38.0 72.3 35.7 77.7 41.36 26.23
Deformable DETR* [50] 54.3 86.0 62.4 96.1 41.07 15.21 38.7 73.6 37.3 87.2 41.08 15.21
PVT-S*[21] 56.2 84.3 62.2 89.9 32.58 12.51 45.6 76.6 47.8 83.6 32.62 12.56
Swin-T*[17] 55.8 89.0 62.3 94.9 36.82 14.58 41.9 75.2 41.3 85.9 36.86 14.64
DAT-T*[28] 47.3 85.3 49.5 92.4 32.05 12.83 41.5 74.3 41.3 85.8 32.09 12.89
BiFormer-S*[18] 49.4 88.6 48.7 95.3 33.62 13.32 45.5 76.8 47.8 85.6 33.66 13.38
MedFormer-T* 52.6 88.8 58.1 95.3 11.09 8.17 42.6 74.6 43.5 86.2 11.13 8.23
MedFormer-S* 53.8 89.3 60.6 93.0 21.22 10.29 43.9 76.1 45.2 84.9 21.27 10.35
MedFormer-B* 56.4 90.1 62.6 94.6 37.59 13.34 45.8 77.1 48.1 84.2 37.63 13.80
@@ 4.2.2 Results
All classification results are summarized in Table 3. It is observed that MedFormer-B achieves the best performance across all three classification datasets. Notably, MedFormer-T achieves comparable performance to some methods with only 3.2M parameters and 0.58G FLOPs. In contrast, HiFuse-Tiny requires significantly more resources but performs well only on the ColonPath dataset. KAT and MG-Trans are effective in reducing computational requirements, but their use of kernel techniques or graph transformers may overlook the extraction of detailed features, thus limiting the improvement in accuracy. The accuracy of MedFormer surpasses that of the compared backbone networks, demonstrating MedFormer’s superior feature representation ability. Additionally, MedFormer with pre-trained weights achieves an accuracy of 100% and 99.91% on the ColonPath and Brain Tumor datasets, respectively, showcasing its immense potential in medical image classification tasks.
It is worth noting that the pretrained version of MedFormer is included solely to demonstrate its potential but is not incorporated into the formal comparisons. To further evaluate the deployment feasibility of these methods in real clinical systems, we measure the throughput and memory usage for each method. The results are shown in Table 3. It can be seen that ViT-S and MIL-VT, although capable of processing images at a high throughput with relatively low memory usage, exhibit lower accuracy. MedFormer-T is suitable for scenarios where speed is prioritized, offering good throughput and low memory usage. MedFormer-B, though slower, excels in accuracy and is ideal for medical image classification tasks requiring high precision. MedFormer-S strikes a good balance between efficiency and accuracy, making it suitable for most clinical applications. These variants ensure that MedFormer can effectively meet the needs of various clinical scenarios.
@@ 4.3.1 Settings
MedFormer is employed as the backbone network in UPerNet [51], and its performance is evaluated against other popular backbone networks on medical image segmentation tasks. Furthermore, comparisons are conducted between MedFormer and several state-of-the-art (SOTA) medical segmentation methods to emphasize its competitiveness. For the ISIC-2018 Segmentation and CVC-ClinicDB datasets, the input image size is set to $\times256$. The Adam optimizer is used with an initial learning rate of 1e41e-4, and training runs for 200 epochs with a cosine annealing learning rate schedule. For the Synapse dataset, the input size is adjusted to $\times224$, the AdamW optimizer is utilized, and training runs for 400 epochs using a cosine annealing learning rate schedule with an initial learning rate of 1e31e-3. To avoid overfitting, data augmentation techniques such as horizontal flipping, vertical flipping, and scaling and rotation are applied. For the key parameters of MedFormer, the dropout rate is set to 0.2, while the region partition factor $S$ is 8 for a resolution of $\times256$ and 7 for $\times224$.
@@ 4.3.2 Results
Table 4 presents the experimental results on the Synapse dataset. In methods without pretraining, MAXFormer achieves leading results, even surpassing some methods that use pretrained weight initialization. In methods with pretrained weights, MedFormer-B* achieves the best average DSC of 84.07%, outperforming all other comparison methods. The result indicates the significant advantage of MedFormer in accurately overlapping the predicted results with the Ground Truth. However, MedFormer does not achieve the highest score in the HD metric, showing a deficiency in learning edge information. We hypothesize that this issue arises because DSSA performs twice sparse selections to focus on a small number of important tokens. Since the number of boundary tokens is much smaller than that of internal organ tokens, and the boundary tokens often have semantic ambiguity, they are overlooked during the selection process. This limitation becomes more evident in multi-organ segmentation datasets with blurry boundaries. Additionally, We observe that while MedFormer-B* achieves the best DSC (84.07%) in the segmentation task, the performance improvement over MAXFormer (83.66%) is relatively small. To verify whether this difference is statistically significant, we performed a paired t-test to compare the performance of the two models on the same dataset. After conducting statistical analysis on the Dice coefficient differences for each sample, we obtain a p-value of 0.048, which is below the common significance level of 0.05. This result indicates that the performance improvement is statistically significant, further supporting the credibility and effectiveness of the proposed method. Furthermore, the segmentation results on the ISIC-2018 Segmentation and CVC-ClinicDB dataset are shown in Table 5. Without using pre-trained weights, MedFormer-B achieves performance comparable to SOTA methods on both datasets. In the comparison of pre-training initialization strategies, MedFormer-B* achieves leading results on the CVC-ClinicDB dataset. Although MedFormer outperforms most competing methods on the ISIC-2018 segmentation dataset, there remains a slight performance gap compared to BRAU-Net++. We consider this result understandable for two reasons: first, the marginal advantage of BRAU-Net++ comes at the cost of significantly higher computational complexity; second, MedFormer is not a dedicated segmentation model and only employs a simple segmentation head—UPerNet. Under the same UPerNet configuration, other backbone networks perform worse than MedFormer. Overall, MedFormer demonstrates robust performance in segmentation tasks, highlighting its potential as a general-purpose backbone network.
@@ 4.4.1 Settings
Since Transformer-based medical image detection methods with available code are scarce, we reproducing several milestone object detection approaches to highlight the competitiveness of MedFormer in detection tasks. Furthermore, we also compare MedFormer with several popular backbone networks under the RetinaNet[48] detection framework to verify its effectiveness. The input resolution is set to $\times256$, and the AdamW optimizer is employed with a cosine annealing learning rate scheduler and a weight decay of 1e41e-4. The initial learning rate is set to 1e41e-4. The parameters for MedFormer are adjusted as follows: the region partition factor $S$ is set to 8, and the drop rate is set to 0.3.
@@ 4.4.2 Results
We use four metrics to evaluate the detection performance of all methods: mean Average Precision ($mAP$), $AP_{50}$, $AP_{75}$, and Recall. The detailed results are presented in Table 6. MedFormer-B achieves the best results on the Kvasir-Seg and Brain Tumor Detection datasets in terms of three metrics: $mAP$, $AP_{50}$ and $AP_{75}$. This proves the strong object detection performance of MedFormer. Furthermore, we observe that the differences in detection results among the comparative methods on the Kvasir-Seg dataset are not significant. However, on the Brain Tumor Detection dataset, dedicated detection methods generally lag behind the backbone networks. We hypothesize that this phenomenon is due to two main reasons: First, the Brain Tumor Detection dataset is multi-class and has nested bounding boxes, which makes accurate localization challenging for the models. Second, the backbone networks involved in the comparison are carefully designed and have stronger feature representation capabilities. All backbone networks are experimented within the RetinaNet framework, replacing the original ResNet encoder, achieving significant improvements. This further confirms that the primary gap exists in the encoder part.
@@
Dual Sparse Selection 88.24 79.44
Table 8: Ablation study on the effect of LCE module on both ISIC-2018 Classification and Synapse Segmentation datasets.
Methods ISIC-2018 Synapse
Top-1 Acc. (%) DSC (%)
MedFormer-S 88.24 79.44
MedFormer-S (w/o LCE) 87.64 78.93
MedFormer-B 89.71 81.75
MedFormer-B (w/o LCE) 88.43 80.54
Table 9: Ablation study of the number of $k_{1}$ and $\lambda$ on ISIC-2018 Classification and CVC-ClinicDB dataset.
$k_{1}$ $\lambda$ ISIC-2018 Classification CVC-ClinicDB
FLOPs (G) Acc (%) FLOPs (G) DSC (%)
1,2,8,49 1/8 2.15 87.05 3.69 87.24
1,4,16,49 1 2.23 87.44 3.79 87.25
1/4 2.18 87.56 3.73 87.64
1/8 2.18 88.24 3.32 88.23
1/16 2.17 85.84 3.31 86.87
2,8,32,49 1/8 2.25 87.56 3.73 87.55
@@
Figure 4: Visualization of attention maps from sample images in the ISIC 2018 Classification dataset (top) and the CVC-ClinicDB segmentation dataset (bottom), illustrating the effects of excluding noise tokens. The pink regions in the segmented samples serve as the ground truth. The red circle represents the query token. It can be seen that pixel-level sparse selection enhances focus on lesion areas, which is particularly important for extracting critical features.
@@ 4.5.1 The effectiveness of DSSA
4.5.2 The effectiveness of LCE
We conduct ablation experiments on the LCE module across different MedFormer variants and tasks, with the results summarized in Table 8. As shown in the results, the LCE module provides a slight but consistent performance improvement across various tasks and model variants. This improvement can be attributed to the module’s use of a $\times5$ depth-wise convolution, which effectively integrates local contextual information while remaining computationally lighter than standard convolutions.
[…]
In this work, we propose two key parameters, $k_{1}$ and $\lambda$, which control the sparse selection at the region-level and pixel-level, respectively. We gradually increase $k_{1}$ at each stage to ensure sufficient feature extraction in smaller regions. Table 9 presents the experimental results for the combination of these parameters. First, based on the analysis of the $k_{1}$ values, we observe that the number of tokens each query can attend to is limited, and introducing too many tokens actually leads to a degradation in performance. Next, we tested three different values of $\lambda$ (with the case of $\lambda=1$ discussed in detail in the next section), and found that the intermediate value of $\lambda=1/8$ yields the best performance on both datasets. We believe this result is due to the model requiring an appropriate number of pixel-level tokens to effectively learn features. Retaining too many pixel-level tokens may introduce excessive noise, distracting the model from key features, while retaining too few may result in insufficient information, limiting the model’s generalization ability. Ultimately, the $k_{1}$ and $\lambda$ hyperparameters we selected achieve the best performance on both datasets, demonstrating their potential generalizability across tasks and datasets.
Figure 5: Visualization of attention maps on ISIC-2018 Classification dataset. For each input, we select two different query positions to observe the region-level tokens they attend to and the pixel-level tokens within those regions. The heatmap values of the pixel-level tokens vary based on their relevance to the query. For this visualization, we use the region-level and pixel-level attention scores extracted from the final MedFormer block of the third stage.
@@ 4.5.4 Effectiveness of Excluding Noise
We design ablation experiments to confirm that pixel-level sparse selection can effectively exclude noise. First, we present quantitative results in Table 9. The accuracy of the model significantly decrease when the selection is not applied ($\lambda=1$). We attribute this drop to the introduction of excessive noise tokens, which prevent the model from extracting relevant features. Second, we visualize the attention map with and without pixel-level sparse selection in Fig. 4. The results clearly show that excluding noise tokens allows the model to focus more on the lesion area. Furthermore, from the segmentation samples, we observed that the attention maps before and after sparse selection exhibit different sensitivities to the boundaries. After sparse selection, the attention converges significantly to the interior of the ground truth boundaries. This suggests that DSSA can, to some extent, adaptively align with the biological boundary features of the lesions.
@@ 4.6.1 Visualization of Attention Map
To intuitively demonstrate the superiority of the proposed DSSA, we visualize the attention maps of DSSA at both the region and pixel levels, as shown in Fig. 5. We observe that DSSA first selects the most relevant regions and then focuses on the tokens most related to the query within those regions. This aligns well with our design philosophy and validates the rationality and effectiveness of our approach. In addition, we visualize the regional attention changes across each layer in MedFormer, as shown in Fig. 6. This visualization shows that as the number of layers increases, our method progressively focuses on relevant regions, demonstrating its ability to refine attention dynamically and align with clinically meaningful areas.
@@ 4.6.2 Visualization of Semantic Segmentation
We visualize the segmentation results of MedFormer, SOTA medical image segmentation methods and popular backbone networks on three datasets. Fig. 7 shows the results on the Synapse dataset, where MedFormer’s segmentation results closely matches the ground truth. In contrast, other methods exhibit issues such as class misidentification and loss of internal details. Fig. 8 presents the results on the ISIC-2018 Segmentation and CVC-ClinicDB datasets. For samples with clear boundaries (Sample 1 and Sample 3), most methods achieve competitive segmentation results. However, for cases with blurred edges or overexposure (Sample 2 and Sample 4), all methods exhibit significant performance degradation and high variability. Notably, MedFormer still maintains leading performance under these challenging conditions, demonstrating strong robustness in complex scenarios. Overall, MedFormer demonstrates superior performance in segmentation tasks due to its unique dual sparse selection attention mechanism, which enables precise extraction of detailed features.
Figure 8: Visualizations of Segmentation using various methods on the CVC-ClinicDB and ISIC-2018 Segmentation datasets. The lesion regions of GT are shown in red, while the segmentation results of methods are shown in pink.